As seen from the window of a space station, a 98.0 kg satellite sailing along at 8.8 m/s collides head-on with a small 318.0 kg asteroid, which was initially at rest. Assume the collision to be elastic. Indicate a change of direction with a negative sign. What is the final velocity of the satellite? What is the final velocity of the asteroid?
Answer #1

final velocity of satellite

v1 = (m1 - m2) u1 / ( m1 + m2)

v1 = ( 98 - 318)* 8.8 / ( 98 + 318)

v1 = - 4.65 m/s

=======

final velocity of asteroid

v2 = 2 m1 u1 / ( m1 + m2)

v2 = 2* 98* 8.8 / ( 98 + 318)

v2 = 4.146 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck